# Physical Thickness Characterization of the FRIB Production Targets

### v1

@@ 1 Introduction
The Facility for Rare Isotope Beams (FRIB), a leading laboratory for rare isotope science, is undergoing a staged ramp-up in primary beam power toward its design goal of 400 kW. Since beginning operations in 2022, FRIB has delivered high-intensity stable ion beams ranging from oxygen to uranium, progressively increasing its operational beam power from 1 kW to 20 kW [1, 2], with current operations approaching 30 kW. This capability supports a broad spectrum of nuclear science experiments, encompassing investigations of nuclear structure, reaction mechanisms, and astrophysical phenomena. To manage the substantial thermal loads generated during rare isotope production, FRIB employs a rotating graphite target configured as a single-slice disc system [3]. Engineered to endure beam powers up to 50 kW, the target mitigates thermal stress by rotating the disc, thereby spreading the energy deposition across a wider surface area. This approach lowers the surface power density and improves thermal efficiency. The current implementation features a graphite disc approximately 30 cm in diameter, rotating at moderate speeds (500 rpm). Plans are underway to increase the rotational speed to 2000 rpm in future operations, which will further suppress the peak temperature and minimize temperature gradients within the irradiated region of the disc. The stable beam, accelerated in the FRIB accelerator section, is delivered to the target with a beam spot size of 0.25 mm (1) in both horizontal and vertical directions. The beam irradiates a 1 cm-wide circular region illustrated in Figure 1 near the rim of the rotating graphite disc. The region maintains a uniform thickness, with the total allowable deviation limited to less than 2 [4]. This tolerance accounts for material inhomogeneity, mechanical deformation, and machining-induced thickness variations. Based on measurements using actual beam irradiation, the combined uncertainty, primarily due to porosity-related inhomogeneity and deformation during 500 rpm rotation, is estimated to be within 1. An additional contribution to the overall uncertainty arises from local variations in physical thickness introduced during the machining process. Currently, the graphite disc is replaced after every ten experimental runs.
@@
To ensure consistent experimental results, accurate characterization of the target thickness is essential. Previously, thickness measurements were conducted at only five discrete positions using a micrometer, introducing significant uncertainty and limited accuracy. To address this limitation, a continuous thickness measurement apparatus based on a laser displacement sensor has been developed. The thickness is measured along the entire circumference with an angular resolution of 0.10 and a radial step size of 1 mm, enabling full areal mapping of the target. Each thickness value represents an average over a circular area with a diameter of 0.5 mm. This method provides a detailed quantification of thickness variations across the entire target surface. We report the characterization of thickness variations in FRIB targets with nominal thicknesses ranging from 0.4 to 5 mm.

### v3

@@ 1 Introduction
The Facility for Rare Isotope Beams (FRIB), a leading laboratory for rare isotope science, is undergoing a staged ramp-up in primary beam power toward its design goal of 400 kW. Since beginning operations in 2022, FRIB has delivered high-intensity stable ion beams ranging from oxygen to uranium, progressively increasing its operational beam power from 1 kW to 20 kW [1, 2], with current operations approaching 30 kW. This capability supports a broad spectrum of nuclear science experiments, encompassing investigations of nuclear structure, reaction mechanisms, and astrophysical phenomena. To manage the substantial thermal loads generated during rare isotope production, FRIB employs a rotating graphite target configured as a single-slice disc system [3]. The target system, shown in Figure 1, consists of a 30 cm-diameter rotating graphite disc with graphite spacers, supported by a YSZ ceramic hub, high-temperature bearings, and an Inconel drive shaft, and enclosed by a copper heat exchanger. The disc assembly provides a total of nine slots to enable future multi-slice disc operation at higher beam power. At present, eight single-slice graphite discs with different thicknesses (1.2, 2.1, 3.5, 5, 8, 10, 15, and 18 mm) are used for secondary beam production. In the near future, sub-mm-thick targets are planned for operation at higher beam power. The operational limit of the graphite disc is governed by sublimation, with a temperature limit of 1900 oC. At a primary beam power of 20 kW and a rotational speed of 500 rpm, the current maximum operational temperature measured is in the range of 800–1200 oC, depending on the target thickness. Eight of these slots are filled with spacers. The graphite grade was selected by considering the minimum grain size (5 m) required for manufacturing a sub-mm-thick disc with a large diameter of 30 cm. In this work, Mersen graphite (grade 2160) and POCO graphite (AXF-5Q) were used. The inner surface of the heat exchanger is coated with a high-emissivity layer ( 0.9) to enhance radiative cooling of the graphite disc. High-temperature bearings and lubricants ( = 280 ) are used to accommodate power ramp-up and high-power operation.
Engineered to endure beam powers up to 50 kW, the target mitigates thermal stress by rotating the disc, thereby spreading the energy deposition across a wider surface area. This approach lowers the surface power density and improves thermal efficiency. The current implementation features a graphite disc approximately 30 cm in diameter, rotating at moderate speeds (500 rpm). Plans are underway to increase the rotational speed to 2000 rpm in future operations, which will further suppress the peak temperature and minimize temperature gradients within the irradiated region of the disc.
@@
The stable beam, accelerated in the FRIB accelerator section, is delivered to the target with a beam spot size of 0.25 mm (1) in both horizontal and vertical directions. The beam irradiates a 1 cm-wide circular region illustrated in Figure 1 near the rim of the rotating graphite disc. The region maintains a uniform thickness, with the total allowable deviation limited to less than 2 [4]. This tolerance accounts for material inhomogeneity, mechanical deformation, and machining-induced thickness variations. Based on measurements using actual beam irradiation, the combined uncertainty, primarily due to porosity-related inhomogeneity and deformation during 500 rpm rotation, is estimated to be within 1. An additional contribution to the overall uncertainty arises from local variations in physical thickness introduced during the machining process. Currently, the graphite disc is replaced after every ten experimental runs. To ensure consistent experimental results, accurate characterization of the target thickness is essential. Previously, thickness measurements were conducted at only five discrete positions using a micrometer, introducing significant uncertainty and limited accuracy. To address this limitation, a continuous thickness measurement apparatus based on a laser displacement sensor, shown in Figure 2, has been developed. The thickness is measured along the entire circumference with an angular resolution of 0.10 and a radial step size of 1 mm, enabling full areal mapping of the target. Each thickness value represents an average over a circular area with a diameter of 0.5 mm. This method provides a detailed quantification of thickness variations across the entire target surface. We report the characterization of thickness variations in FRIB targets with nominal thicknesses ranging from 0.4 to 5 mm.